# Dynamics of elastic wires: Preserving Area Without Nonlocality

### v1

@@ 6.3. Proof of full convergence
For the proof of full convergence, we use an argumentation similar to [11, Section 5], but due to the $H^{-1}$-structure of the flow, applying the Łojasiewicz inequality is more challenging and time-dependent norms have to be handled. In [19], the authors have to face similar difficulties. We note that in [44], an elegant way to shorten the proof of convergence in [11, Section 5] is shown, but it seems that this cannot be transferred to $H^{-1}$-flows.
@@
We assume that the inequality is strict. If not, there exists $\bar{t}\in[0,\infty)$ such that $\mathcal{E}_{\lambda}(\gamma(\bar{t}))=\mathcal{E}_{\lambda}(\gamma_{\infty})$. But due to 3.4, this means $\mathcal{E}_{\lambda}(\gamma(t))=\mathcal{E}_{\lambda}(\gamma_{\infty})$ for all $t\geq\bar{t}$, so $\partial_{t}^{\perp}\gamma\equiv 0$ on $[\bar{t},\infty)$ by (3.13). Thus, only the tangential velocity changes on on $[\bar{t},\infty)$ and the claimed convergence follows trivially.
@@
for all vector fields $\alpha$ normal to $\gamma_{\infty}$. The latter equation makes sense, because
@@
becomes small for $\|\zeta-\gamma_{\infty}\|_{C^{1}}$ small.
@@
By 4.1 and 4.2, there exists a a geometrically unique smooth solution $\bar{\gamma}$ of (6.44) that can be written as $\bar{\gamma}(t,x)=\gamma_{\infty}(x)+\varphi(t)\nu_{\infty}(x)$ for some smooth scalar function $\varphi$. Let $0<T^{\prime}\leq\infty$ the maximal existence time of this solution. The geometric uniqueness implies that on $[t_{i},T^{\prime})$, $\bar{\gamma}$ is a smooth reparametrization of $\tilde{\gamma}-p(t_{i})$. It follows that for proving 6.6, it suffices to prove that $T^{\prime}=\infty$ and that $\bar{\gamma}$ converges to $\gamma_{\infty}$ as $t\to\infty$.
Step 5: $T$ and $T^{\prime\prime}$. Let $T\in(0,T^{\prime}]$ be maximal such that
@@
Due to (6.43), such a $T$ exists. Our goal is to show $T=T^{\prime}=\infty$. For this, we assume by contradiction that either $0<T<T^{\prime}=\infty$ or $0<T\leq T^{\prime}<\infty$. With this assumption, we define the finite time $T^{\prime\prime}:=\min\{T^{\prime},T+1\}$. Analogously to [11, p. 2190 et seq.] it can be shown that
(6.46) $\sup_{[0,T)}\|\varphi(t)\|_{C^{6,\alpha}}\leq C(\delta,\tilde{r},T^{\prime\prime})\quad\text{ and hence }\quad\sup_{[0,T)}\|\bar{\gamma}(t)-\gamma_{\infty}\|_{C^{6,\alpha}}\leq C(\delta,\tilde{r},T^{\prime\prime})$
@@
Step 6: Application of the Łojasiewicz-Simon inequality. We define
@@
Step 7: Time-independent norm. With the same partial integration as in 3.6, we see that $\nabla_{s}^{2}\nabla_{L^{2}}\mathcal{E}_{\lambda}(\bar{\gamma})\in H_{\bar{\gamma}}$ and hence $\partial_{t}^{\perp}\bar{\gamma}\in H_{\bar{\gamma}}$. So 2.7 implies that
@@
where $Y^{\perp_{\infty}}=\langle Y,\nu_{\infty}\rangle\nu_{\infty}$ and $W^{1,2,{\perp_{\infty}}}=W^{1,2}\cap\{Y\perp\gamma_{\infty}\}$. For the first term in the last inequality, we used a computation as in (2.34) and for the second term we used $\partial_{t}\bar{\gamma}=\langle\partial_{t}\bar{\gamma},\nu_{\infty}\rangle\nu_{\infty}$. Since $1=|\nu_{\infty}|^{2}=|\langle\nu_{\infty},\bar{\nu}\rangle|^{2}+|\langle\nu_{\infty},\partial_{s}\bar{\gamma}\rangle|^{2}$ and $|\langle\nu_{\infty},\bar{\nu}\rangle|\geq\frac{1}{2}$ on $(0,T)$ by (6.39), we conclude that
@@
Step 8: Interpolation. Since $N$ is normal to $\gamma_{\infty}$ and since $\mathcal{A}(\bar{\gamma}(t))=\mathcal{A}(\gamma_{\infty})=A_{0}$ implies $\int_{\mathbb{S}^{1}}\langle N(t),\nu_{\infty}\rangle\,\mathrm{d}s_{\gamma_{\infty}}=0$, we have $\bar{\gamma}(t)-\gamma_{\infty}=N(t)\in H_{\gamma_{\infty}}$. Thus, 2.6 and 2.7 yield
@@
where we used (6.65) and $\|\bar{\gamma}(t)-\gamma_{\infty}\|_{H_{\gamma_{\infty}}}\leq C\|\bar{\gamma}(t)-\gamma_{\infty}\|_{W^{6,2}}<\delta$, see (6.45). With [32, Prop. 1.1.3], we have
@@
(6.73) $\leq C(\tilde{r},\delta,T^{\prime\prime})\|\bar{\gamma}(t)-\gamma_{\infty}\|_{L^{2}(\,\mathrm{d}s_{\gamma_{\infty}})}^{\beta}$
@@
Step 9: $T=T^{\prime}=\infty$. Choosing $\varepsilon<\big{(}\frac{\min\{\delta,\tilde{r}\}}{2C}\big{)}^{\frac{2}{\vartheta\beta}}$, we obtain from (6.74)
@@
Step 10: Convergence. With (6.76) and (6.62), we obtain $\|\partial_{t}\bar{\gamma}\|_{\star}\in L^{1}(\mathbb{R}_{+})$. From
@@ 6.4. Stationary solutions
If $c_{1}=0$ and $\lambda>0$, we obtain the classical elasticae equation (1.3), for which the solutions are classified. So in a certain sense, the constant $c_{1}$ measures the deviation from the stationary solution to an elastica. Remember that the only closed elasticae are multifold coverings of the circle and of the figure eight elastica, see [38, Lemma 5.4].
@@
with penalized area ($\lambda>0$ given) and penalized length ($c_{1}>0$ given). In [45], a complete classification of the critical points corresponding to closed curves is presented. Similarly, [46] explicitly provides the solutions to (6.78) in a format comparable to [45]. The authors state a condition under which the solution describes a closed curve and investigate when the curve is simple. By [9, Lemma 5.4], minimizing (6.79) is equivalent to minimize $\mathcal{E}_{\lambda}$ under the constraint of fixed enclosed area. The latter problem is studied in [48] and [49] for $\omega=1$. In particular, we note that from the mentioned articles, it follows that for $c_{1}\neq 0$, there exist solutions of (6.78) that do not correspond to a multifold covered circle or a multifold covered figure eight. Concrete examples can be found in [45], [46] or [49]. We note that the references mentioned are by no means complete. For further references, see also the citations within the mentioned articles.
@@ 6.5. Blow-up in infinite time
One might think that for $A_{0}\neq 0$ and $\lambda=0$, the length of a curve can be controlled by a bound on the elastic energy. We expect that this is not the case. We have in mind a curve shaped like a dumbbell, where we can make the bar longer and thinner, allowing us to increase the length without increasing the elastic energy or the area. This means that also for $A_{0}\neq 0$, in general, we do not expect the length to remain bounded, which would imply convergence for $\lambda=0$.
@@ A.1. A crucial reparametrization argument
For the convenience of the reader, we state the following reparametrization argument, which can be found with a detailed proof in [11, Lemma 4.1 and Remark 4.2]. In addition to the proof of short time existence (4.1), it is also used in the proof of the constrained Łojasiewicz–Simon gradient inequality (6.5).
@@
With the notations in the preceding lemma, it can be shown that smoothness of the reference curve $\hat{\gamma}$ and smoothness of $X\in B_{r}(0)$ yields smoothness of the diffeomorphism $\Phi$ and the normal variation $Y$. For details on this, see [11, Proof of Lemma 4.1].
@@ A.2. Parabolic Hölder spaces and the linear problem
For the proof of short time existence, we work with parabolic Hölder spaces of order $6$, cf. [15, Section II.3.1] or [26].
@@
First, assume that $a_{6}\equiv 1$ and $a_{q}\equiv 0$ for $q=0,1,\dots,5$. Like in [23, Section 3] one proves that in this case, a weak solution to (A.6) exists. Localizing the problem and using [15, Theorem VI.21] one obtains Hölder regularity and (A.7). As in [3, Section 3.1.4], the general linear problem (A.6) is solved with the method of continuity. ∎
@@ A.3. The nonlinear problem
In this subsection, we proceed as in [21, Section 2.5] and [12, Section 3.3]. For a smooth immersed reference curve $\gamma_{\ast}$ with normal $\nu_{\ast}$, we define the open set
@@
Analogously as in [21, Lemma 2.5.8], one shows that $f_{\varepsilon}\in H^{\frac{l}{6},l}\left([0,T]\times{\mathbb{S}^{1}}\right)$ with norm bounded independently of $\varepsilon$. Here, we need the assumption $l<1$. The crucial point is estimating the Hölder seminorm in time. We have
@@
Again, the constant $C$ does not depend on $\varepsilon$, which allows us to show the claim. Moreover, one proves with standard arguments (see for example [22, Section 6.8]) and the Arzelà–Ascoli theorem, that for $10<k<l<$ 1, $H^{\frac{l}{6},l}\left([0,T]\times{\mathbb{S}^{1}}\right)$ is compactly embedded in $H^{\frac{k}{6},k}\left([0,T]\times{\mathbb{S}^{1}}\right)$. It follows that there exists a sequence $\varepsilon_{n}\to 0$, $n\to\infty$, and $\hat{f}\in H^{\frac{l}{6},l}\left([0,T]\times{\mathbb{S}^{1}}\right)$ such that
@@
For the terms coming from the Hölder seminorm in space, interpolation as in [32, Prop. 1.1.3] yields

### v2

@@ 6.3. Proof of full convergence
For the proof of full convergence, we use an argumentation similar to [13, Section 5], but due to the $H^{-1}$-structure of the flow, applying the Łojasiewicz inequality is more challenging and time-dependent norms have to be handled. In [21], the authors have to face similar difficulties. We note that in [46], an elegant way to shorten the proof of convergence in [13, Section 5] is shown, but it seems that this cannot be transferred to $H^{-1}$-flows.
@@
We assume that the inequality is strict. If not, there exists $\bar{t}\in[0,\infty)$ such that $\mathcal{E}_{\lambda}(\gamma(\bar{t}))=\mathcal{E}_{\lambda}(\gamma_{\infty})$. But due to 3.3, this means $\mathcal{E}_{\lambda}(\gamma(t))=\mathcal{E}_{\lambda}(\gamma_{\infty})$ for all $t\geq\bar{t}$, so $\partial_{t}^{\perp}\gamma\equiv 0$ on $[\bar{t},\infty)$ by (3.12). Thus, only the tangential velocity changes on on $[\bar{t},\infty)$ and the claimed convergence follows trivially.
@@
for all vector fields $\alpha$ normal to $\gamma_{\infty}$. The latter inequality can be achieved by
@@
becoming small for small distance $\|\zeta-\gamma_{\infty}\|_{C^{1}}$.
@@
By 4.1 and 4.2, there exists a a geometrically unique smooth solution $\bar{\gamma}$ of (6.44) that can be written as $\bar{\gamma}(t,x)=\gamma_{\infty}(x)+\varphi(t,x)\nu_{\infty}(x)$ for some smooth scalar function $\varphi$. Let $0<T^{\prime}\leq\infty$ the maximal existence time of this solution. The geometric uniqueness implies that on $[t_{i},T^{\prime})$, $\bar{\gamma}$ is a smooth reparametrization of $\tilde{\gamma}-p(t_{i})$. It follows that for proving 6.6, it suffices to prove that $T^{\prime}=\infty$ and that $\bar{\gamma}$ converges to $\gamma_{\infty}$ as $t\to\infty$.
Step 4: Let $T\in(0,T^{\prime}]$ be maximal such that
@@
Due to (6.43), such a $T$ exists. Our goal is to show $T=T^{\prime}=\infty$. By parabolic Schauder estimates (as in [8, Section 4]), it can be shown that
(6.46) $\sup_{[0,T)}\|\bar{\gamma}(t)-\gamma_{\infty}\|_{C^{6,\alpha}}\leq C(\delta,\tilde{r},\gamma_{0})$
@@
Step 5: Application of the Łojasiewicz-Simon inequality. We define
@@
Step 6: Time-independent norm. With the same partial integration as in 3.5, we see that $\nabla_{s}^{2}\nabla_{L^{2}}\mathcal{E}_{\lambda}(\bar{\gamma})\in H_{\bar{\gamma}}$ and hence $\partial_{t}^{\perp}\bar{\gamma}\in H_{\bar{\gamma}}$. So 2.7 implies that
@@
where $Y^{\perp_{\infty}}=\langle Y,\nu_{\infty}\rangle\nu_{\infty}$ and $W^{1,2,{\perp_{\infty}}}=W^{1,2}\cap\{Y=Y^{\perp_{\infty}}\}$. For the first term in the last inequality, we used a computation as in (2.34) and for the second term we used $\partial_{t}\bar{\gamma}=\langle\partial_{t}\bar{\gamma},\nu_{\infty}\rangle\nu_{\infty}$. Since $1=|\nu_{\infty}|^{2}=|\langle\nu_{\infty},\bar{\nu}\rangle|^{2}+|\langle\nu_{\infty},\partial_{s}\bar{\gamma}\rangle|^{2}$ and $|\langle\nu_{\infty},\bar{\nu}\rangle|\geq\frac{1}{2}$ on $(0,T)$ by (6.39), we conclude that
@@
Step 7: Interpolation. Since $N$ is normal to $\gamma_{\infty}$ and since $\mathcal{A}(\bar{\gamma}(t))=\mathcal{A}(\gamma_{\infty})=A_{0}$ implies $\int_{\mathbb{S}^{1}}\langle N(t),\nu_{\infty}\rangle\,\mathrm{d}s_{\gamma_{\infty}}=0$, we have $\bar{\gamma}(t)-\gamma_{\infty}=N(t)\in H_{\gamma_{\infty}}$. Thus, 2.6 and 2.7 yield
@@
where we used (6.65) and $\|\bar{\gamma}(t)-\gamma_{\infty}\|_{H_{\gamma_{\infty}}}\leq C\|\bar{\gamma}(t)-\gamma_{\infty}\|_{W^{6,2}}<\delta$, see (6.45). With [34, Prop. 1.1.3], we have
@@
(6.73) $\leq C(\tilde{r},\delta,\gamma_{0})\|\bar{\gamma}(t)-\gamma_{\infty}\|_{L^{2}(\,\mathrm{d}s_{\gamma_{\infty}})}^{\beta}$
@@
Step 8: $T=T^{\prime}=\infty$. Choosing $\varepsilon<\big{(}\frac{\min\{\delta,\tilde{r}\}}{2C}\big{)}^{\frac{2}{\vartheta\beta}}$, we obtain from (6.74)
@@
Step 9: Convergence. With (6.76) and (6.62), we obtain $\|\partial_{t}\bar{\gamma}\|_{\star}\in L^{1}(\mathbb{R}_{+})$. From
@@ 6.4. Stationary solutions
If $c_{1}=0$ and $\lambda>0$, we obtain the classical elasticae equation (1.3), for which the solutions are classified. So in a certain sense, the constant $c_{1}$ measures the deviation from the stationary solution to an elastica. Remember that the only closed elasticae are multifold coverings of the circle and of the figure eight elastica, see [40, Lemma 5.4].
@@
with penalized area ($\lambda>0$ given) and penalized length ($c_{1}>0$ given). In [47], a complete classification of the critical points corresponding to closed curves is presented. Similarly, [48] explicitly provides the solutions to (6.78) in a format comparable to [47]. The authors state a condition under which the solution describes a closed curve and investigate when the curve is simple. By [11, Lemma 5.4], minimizing (6.79) is equivalent to minimize $\mathcal{E}_{\lambda}$ under the constraint of fixed enclosed area. The latter problem is studied in [50] and [51] for $\omega=1$. In particular, we note that from the mentioned articles, it follows that for $c_{1}\neq 0$, there exist solutions of (6.78) that do not correspond to a multifold covered circle or a multifold covered figure eight. Concrete examples can be found in [47], [48] or [51]. We note that the references mentioned are by no means complete. For further references, see also the citations within the mentioned articles.
@@ 6.5. Blow-up in infinite time
We note that also for $A_{0}\neq 0$ and $\lambda=0$, the length of the curve is not controlled by a bound on the elastic energy in general, see for example [6].
@@ A.1. A crucial reparametrization argument
For the convenience of the reader, we state the following reparametrization argument, which can be found with a detailed proof in [13, Lemma 4.1 and Remark 4.2]. In addition to the proof of short time existence (4.1), it is also used in the proof of the constrained Łojasiewicz–Simon gradient inequality (6.5).
@@
With the notations in the preceding lemma, it can be shown that smoothness of the reference curve $\hat{\gamma}$ and smoothness of $X\in B_{r}(0)$ yields smoothness of the diffeomorphism $\Phi$ and the normal variation $Y$. For details on this, see [13, Proof of Lemma 4.1].
@@ A.2. Parabolic Hölder spaces and the linear problem
For the proof of short time existence, we work with parabolic Hölder spaces of order $6$, cf. [17, Section II.3.1] or [28].
@@
First, assume that $a_{6}\equiv 1$ and $a_{q}\equiv 0$ for $q=0,1,\dots,5$. Like in [25, Section 3] one proves that in this case, a weak solution to (A.6) exists. Localizing the problem and using [17, Theorem VI.21] one obtains Hölder regularity and (A.7). As in [3, Section 3.1.4], the general linear problem (A.6) is solved with the method of continuity. ∎
@@ A.3. The nonlinear problem
In this subsection, we proceed as in [23, Section 2.5] and [14, Section 3.3]. For a smooth immersed reference curve $\gamma_{\ast}$ with normal $\nu_{\ast}$, we define the open set
@@
Analogously as in [23, Lemma 2.5.8], one shows that $f_{\varepsilon}\in H^{\frac{l}{6},l}\left([0,T]\times{\mathbb{S}^{1}}\right)$ with norm bounded independently of $\varepsilon$. Here, we need the assumption $l<1$. The crucial point is estimating the Hölder seminorm in time. We have
@@
Again, the constant $C$ does not depend on $\varepsilon$, which allows us to show the claim. Moreover, one proves with standard arguments (see for example [24, Section 6.8]) and the Arzelà–Ascoli theorem, that for $10<k<l<$ 1, $H^{\frac{l}{6},l}\left([0,T]\times{\mathbb{S}^{1}}\right)$ is compactly embedded in $H^{\frac{k}{6},k}\left([0,T]\times{\mathbb{S}^{1}}\right)$. It follows that there exists a sequence $\varepsilon_{n}\to 0$, $n\to\infty$, and $\hat{f}\in H^{\frac{l}{6},l}\left([0,T]\times{\mathbb{S}^{1}}\right)$ such that
@@
For the terms coming from the Hölder seminorm in space, interpolation as in [34, Prop. 1.1.3] yields